OpenStudy (anonymous):

HELP PLEASE WILL GIVE A MEDAL ") A bag contains 13 dark chocolates, 16 white chocolates, and 11 milk chocolates. Maggie’s class teacher, Jill, is very strict, and the probability that she offers her students chocolates is 0.4. If Jill offers Maggie a chocolate from the bag, what is the probability that she randomly picks either a white chocolate or a milk chocolate?

OpenStudy (anonymous):

A. 11/10 B. 27/40 C. 4/7 D. 4/10 E. 13/40

OpenStudy (anonymous):

I got E is that correct?

OpenStudy (shinalcantara):

i think it's B ----------- Probability of picking either white or milk chocolate \[\frac{ _{16}C_1 }{ 40 } + \frac{ _{11}C_1 }{ 40 } = \frac{ 27 }{ 40 }\]

OpenStudy (anonymous):

thanks